OpenStudy (anonymous):

what is f'(x) = [2x^2tan(x)] / [sec(x)] ?

OpenStudy (anonymous):

i know you have to use the quotient rule and so far I got f'(x) = (4x)(tan(x)) - (sec^2(x))(2x^2) / [sec^2(x)]

OpenStudy (anonymous):

but how do you simplify this?

OpenStudy (anonymous):

Is it \(f(x) = \frac{2x^2 \cdot \tan(x)}{\sec(x)}\)? and you want to find \(f'(x)\) now?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Ok, well in that case, what is \(\sec(x)\)?

OpenStudy (anonymous):

1/cos(x)

OpenStudy (anonymous):

Right, so we can say: $$ f(x) = \frac{2x^2 \tan(x)}{\sec(x)} = = \frac{2x^2 \tan(x)}{\frac{1}{\cos(x)}} $$and we can multiply both the numerator and denominator by \(\cos(x)\) and get: $$ f(x) = \frac{2x^2 \cdot \tan(x) \cdot \cos(x)}{\frac{1}{\cos(x)} \cdot \cos(x)} = 2x^2 \cdot \tan(x) \cdot \cos(x) $$

OpenStudy (anonymous):

I wrote accidently == instead of = Now, what is \(\tan(x)\)?

OpenStudy (anonymous):

sin(x)/cos(x)

OpenStudy (anonymous):

Right, so again: $$ f(x) = 2x^2 \cdot \tan(x) \cdot \cos(x) = 2x^2 \cdot \frac{\sin(x)}{\cos(x)} \cdot \cos(x) = 2x^2 \cdot \sin(x) $$

OpenStudy (anonymous):

Now it should be much easier to take the derivative using the product rule

OpenStudy (anonymous):

So what do you get?

OpenStudy (anonymous):

yeah i got 4xsin(x) + 2x^2cos(x) and then i simplified it to 2x(2sin(x) + xcos(x))

OpenStudy (anonymous):

okay wow thank you so much! when I first saw this problem I was like how in the world do you solve this

OpenStudy (anonymous):

lol, you're welcome =)

OpenStudy (anonymous):

so in order to find f'(5) do you just plug in 5 to final answer?

OpenStudy (anonymous):

For \(x\) yes

OpenStudy (anonymous):

That is the power of derivative, instead of finding a slope in a single point you generate a function for all points

OpenStudy (anonymous):

Also, in worst case, that you can't simplify it like we did here, try to do it in parts. First call the numerator \(u\) or something, and use the quotient rule and then calculate \(u'\) separately and plug in wherever you had \(u'\) in the first step. So in our case we could say: $$ f(x) = \frac{2x^2 \cdot \tan(x)}{\sec(x)} \\ u = 2x^2 \cdot \tan(x) \qquad u'=(4x \cdot \tan(x)) + (2x^2 \cdot \sec^2(x)) \\ f(x) = \frac{u}{\sec(x)}\\ f'(x) = \frac{(u' \cdot \sec(x)) - (u \cdot \tan(x) \cdot \sec(x))}{\sec^2(x)} $$And now we can plug in $$ f'(x) = \frac{(4x \cdot \tan(x) + 2x^2 \cdot \sec^2(x)) \cdot \sec(x) - 2x^2 \cdot \tan(x) \cdot \tan(x) \cdot \sec(x)}{\sec^2(x)} $$And simplify.. which in our case we could from the beginning.. This is of course very ugly.. but if you have no choice then this is the approach I'd take.

OpenStudy (anonymous):

i don't think i'm getting the right answer. i keep getting this weird decimal .0177789598

OpenStudy (anonymous):

Well \(\sin(5)\) and \(\cos(5)\) will not give you nice values.. Are you sure you're expecting one for \(x=5\)?

OpenStudy (anonymous):

well the question says what is f'(5) using the f(x) that I wrote before. it's part of the same question. first they tell you to find f'(x) and then f'(5)

OpenStudy (anonymous):

when i plugged my answer into the computer it marked it wrong

OpenStudy (anonymous):

The result it gives here is -4.99538 not 0.0177...

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

make sure that when you calculate sin and cos you do it by radians

OpenStudy (anonymous):

ohhh i had it in degree

OpenStudy (anonymous):

the derivative \((\sin(x))' = \cos(x)\) is defined that way only when you're working in radians. It can be shown from the definition of the derivative, but if you didn't get there yet then just remember it =)

OpenStudy (anonymous):

i'm still not getting in right. pretty sure i'm plugging it wrong into my calculator though

OpenStudy (anonymous):

*it

OpenStudy (anonymous):

did you try rounding to -5?

OpenStudy (anonymous):

oh no my number's not even close to 5. you do plug in 5 for x in this equation right? : f′(x)=(4x⋅tan(x)+2x^2⋅sec2^(x))⋅sec(x)−2x^2⋅tan(x)⋅tan(x)⋅sec(x)sec^2(x)

OpenStudy (anonymous):

why are you using this terrible form? use the short one we found $$f'(x) = 2x(2\sin(x) + x\cos(x))$$ if you plug \(x=5\) this becomes $$f'(5) = 2 \cdot 5(2\sin(5) + 5\cos(5)) = 20 \sin(5) + 50 \cos(5) = ?$$

OpenStudy (anonymous):

oh wow i'm dumb. yeah you get the right answer this way

OpenStudy (anonymous):

thanks so much once again!

OpenStudy (anonymous):

sure no problem =)

OpenStudy (anonymous):

so are you always supposed to simplify as much as you can before finding the derivative to make your life easier or did that only work in this case?